# Critical LAN and Score Tests for Mixed Fractional Models under High-Frequency Observation at $H=3/4$

### 24 Mar 2026 v1

@@ 7 Simulation study
We then evaluate the mfBm score components at $(\widehat{\sigma}_{0},3/4)$ and form the plug-in efficient statistic
where
$\widehat{\Xi}_{\sigma,n}^{\mathrm{m}}=\frac{S_{\sigma,n}^{\mathrm{m}}(\widehat{\sigma}_{0},3/4)}{\sqrt{n\Delta_{n}L_{n}}},\qquad\widehat{\Xi}_{H,n}^{\mathrm{m}}=\frac{R_{H,n}^{\mathrm{m}}(\widehat{\sigma}_{0},3/4)}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}.$
The one-sided test rejects for large positive values of $\widehat{T}_{n}^{\mathrm{m}}$.
We present two numerical experiments.
Experiment I: Fixed design finite sample performance.
This experiment keeps the sampling design fixed and asks whether the statistic defined above already moves in the correct direction when the true Hurst index crosses from the critical point into the supercritical region. We take
@@
and use $200$ Monte Carlo replications in each regime. We compare
$H=3/4\qquad\text{and}\qquad H=0.90.$
The first is the critical null hypothesis, while the second is a clearly supercritical alternative. For each replication we compute $\widehat{\sigma}_{0}$, then $\widehat{T}_{n}^{\mathrm{m}}$, and finally the empirical rejection frequencies at the one-sided $10\%$ and $5\%$ Gaussian critical values.
Experiment II: Critical null hypothesis behavior along an asymptotic sequence.
The second experiment keeps the true model at the critical point
@@
and examines how the plug-in fit and the resulting statistic behave along a simple asymptotic sequence. To approximate the LAN setting with $\Delta_{n}\downarrow 0$ and $n\Delta_{n}\to\infty$, we take
$\Delta_{n}=n^{-1/2},\qquad n\in\{64,128,256\},$
again with $200$ Monte Carlo replications for each $n$. The goal is descriptive: we track the numerical stability of $\widehat{\sigma}_{0}$ and compare the empirical behavior of $\widehat{T}_{n}^{\mathrm{m}}$ with its $\mathcal{N}(0,1)$ benchmark as the sample size increases.
Table 1: Experiment I: feasible one-sided mfBm test at a fixed design
@@
$0.75$ 0.003-0.003 $0.532$ $0.025$ $0.005$ $3.034$ $0.377$
$0.90$ $1.461$ $1.726$ $0.445$ $0.380$ $1.519$ $0.456$
Table 2: Experiment II: critical-null hypothesis behavior of the plug-in mfBm statistic along an asymptotic sequence
@@
$64$ $0.1250$ $0.006$ $0.751$ $0.045$ $2.948$ $0.358$
$128$ $0.0884$ $0.033$ $0.613$ $0.035$ $2.961$ $0.308$
$256$ $0.0625$ 0.047-0.047 $0.491$ $0.005$ $2.988$ $0.216$
The simulations show that the feasible mfBm statistic is numerically implementable once the restricted estimator $\widehat{\sigma}_{0}$ is obtained under $H=3/4$. Under the critical null hypothesis it is centered near zero but remains conservative in finite samples, suggesting slow convergence to the asymptotic normal limit. Under the supercritical alternative it shifts to the right and clearly improves rejection. Thus the current experiments support practical implementability and correct directional behavior, while a formal plug-in justification is left for future work.
Appendix A Critical low frequency reductions and explicit constants
@@ A.4 Higher-order critical profiles and LAN remainders
$\left\lVert N_{\beta,n}^{\bullet}\right\rVert_{\mathrm{op}}=O\bigl(L_{n}^{m(\beta)}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(N_{\beta,n}^{\bullet}\bigr)^{2}\Bigr)=O\bigl(n\Delta_{n}L_{n}^{2m(\beta)+1}\bigr),\qquad|\beta|=2,3.$ (56)
In particular, every additional $H$-derivative costs exactly one extra logarithmic factor at the operator level and two at the trace level.
@@
$\left\lVert\mathcal{R}_{2,n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{2,n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$ (58)
$\left\lVert\mathcal{R}_{2,n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{2,n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n})^{-1}\bigr),$ (59)
$\left\lVert\mathcal{R}_{3,n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-3/2}L_{n}^{-3/2}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{3,n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n})^{-2}L_{n}^{-2}\bigr),$ (60)
$\left\lVert\mathcal{R}_{3,n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-3/2}\bigr),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{3,n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)$ $=O\bigl((n\Delta_{n})^{-2}\bigr).$ (61)
@@
and proposition 21 shows that each of them is of operator order $(n\Delta_{n}L_{n})^{-1}$ and squared-trace order $(n\Delta_{n}L_{n})^{-1}$. This yields (58). For mfOU the additional $(\sigma,\alpha)$, $(H,\alpha)$ and $(\alpha,\alpha)$ terms are even smaller, except for $(\alpha,\alpha)$ whose size is exactly $(n\Delta_{n})^{-1}$; hence (59) follows.
For the cubic remainder, every monomial contains three local shifts and a third derivative. In mfBm the worst case is any monomial with total $m(\beta)=3$ after multiplying by $(\delta_{H,n})^{3}$, and all such contributions have the same size
$(n\Delta_{n})^{-3/2}L_{n}^{-9/2}\cdot L_{n}^{3}=(n\Delta_{n})^{-3/2}L_{n}^{-3/2}$

### 02 Apr 2026 v2

@@ 7 Monte Carlo diagnostics for the feasible boundary statistic
In every replication we simulate the mfBm increment vector under the relevant data-generating law, re-estimate $\widehat{\sigma}_{0}$ under $H=3/4$, and then evaluate $\widehat{T}_{n}^{\mathrm{m}}$ exactly as in Section 6. Large positive values support $H>3/4$.
The first experiment studies directional power at a fixed high-frequency design,
@@
over the grid
$H\in\{0.75,0.78,0.80,0.85,0.90\}.$
The second experiment keeps the model at the boundary,
@@
and follows the asymptotic sequence
$\Delta_{n}=n^{-1/2},\qquad n\in\{64,128,256,512\},$
again with 1,0001{,}000 replications at each design point.
Table 1: Experiment I: power curve of the feasible one-sided mfBm statistic at a fixed design
@@
$0.75$ $0.013$ $0.528$ $0.022$ $0.009$ $2.982$ $0.397$
$0.78$ $0.190$ $0.761$ $0.074$ $0.038$ $2.540$ $0.392$
$0.80$ $0.386$ $0.854$ $0.134$ $0.080$ $2.339$ $0.392$
$0.85$ $0.854$ $1.375$ $0.321$ $0.245$ $1.839$ $0.444$
$0.90$ $1.422$ $1.846$ $0.474$ $0.394$ $1.472$ $0.494$
Table 2: Experiment II: critical-null behavior of the feasible mfBm statistic along an asymptotic sequence
@@
$64$ $0.1250$ $0.018$ $0.688$ $0.044$ $0.023$ $2.990$ $0.378$
$128$ $0.0884$ 0.013-0.013 $0.593$ $0.029$ $0.009$ $2.973$ $0.283$
$256$ $0.0625$ $0.011$ $0.549$ $0.026$ $0.006$ $2.995$ $0.211$
$512$ $0.0442$ 0.034-0.034 $0.511$ $0.007$ $0.002$ $2.995$ $0.152$
Table 1 shows that the feasible statistic is directionally informative in the way the theory predicts. Under the boundary law it remains essentially centered and heavily conservative, but its mean and rejection frequency both increase as the true Hurst index moves deeper into the supercritical region. The power gain is gradual near $H=0.78$ and $H=0.80$, then becomes much more visible at $H=0.85$ and $H=0.90$.
Table 2 shows equally clearly that the conservative null behavior is not a small-sample accident. Even as $n$ increases from $64$ to $512$, the empirical standard deviation of $\widehat{T}_{n}^{\mathrm{m}}$ stays far below the asymptotic unit benchmark and the rejection rates at nominal $10\%$ and $5\%$ levels remain well below target. The statistic is therefore numerically stable and already useful as a one-sided diagnostic, but first-order Gaussian critical values still under-reject substantially at realistic sample sizes.
Figure 1 and Figure 2 summarize the same two messages graphically. The first is that the feasible score has a genuine power curve once the data move above the boundary. The second is that the critical logarithmic normalization makes convergence slow, so we should eventually supplement the asymptotic critical values with a finite-sample correction such as a null-based bootstrap.
Figure 1: Empirical rejection rates of the feasible mfBm boundary score across the supercritical power grid in Experiment I.
Figure 2: Empirical standard deviation of the feasible mfBm boundary score under the critical null across the asymptotic sequence in Experiment II.
The Monte Carlo section is not definitive, but the upgraded simulations already show that the feasible mfBm statistic is stable, its power increases as expected, and its main shortcoming is conservative null calibration, not numerical instability.
8 Empirical implementation of the exact feasible boundary score
We apply the exact feasible mfBm statistic to one-minute SPY data from 2008-01-22 to 2021-05-06. Because overnight gaps break the equally spaced increment scheme, we work day by day on the regular session only, keeping observations from 07:30 to 13:59 and discarding incomplete days. This leaves 2,6212{,}621 trading days, each with $390$ prices and therefore $n=389$ close-to-close log returns.
For each day we first normalize the intraday return vector so that its empirical diffusive variance matches the Brownian benchmark $\Delta_{n}=1/389$. Under the boundary restriction $H=3/4$ we then estimate $\sigma$ by the exact Gaussian likelihood,
$\widehat{\sigma}_{0}:=\arg\max_{\sigma>0}\,\ell_{n}^{\mathrm{m}}(\sigma,3/4),$
and evaluate
$\widehat{\Xi}_{\sigma,n}^{\mathrm{m}}:=\frac{S_{\sigma,n}^{\mathrm{m}}(\widehat{\sigma}_{0},3/4)}{\sqrt{n\Delta_{n}L_{n}}},\qquad\widehat{\Xi}_{H,n}^{\mathrm{m}}:=\frac{R_{H,n}^{\mathrm{m}}(\widehat{\sigma}_{0},3/4)}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}.$
The empirical boundary statistic is therefore
$\widehat{T}_{n}^{\mathrm{m}}:=\frac{8}{\sqrt{3}\,\widehat{\sigma}_{0}^{2}}\left(\widehat{\Xi}_{H,n}^{\mathrm{m}}-\frac{\widehat{\sigma}_{0}}{2}\widehat{\Xi}_{\sigma,n}^{\mathrm{m}}\right).$ (42)
Large positive values support the one-sided alternative $H>3/4$.
Table 3 summarizes the daily statistic over the full sample. The center of the distribution stays close to zero, the empirical dispersion is slightly below one, and the rejection shares are modest: $10.4\%$ at the nominal $10\%$ threshold and $6.5\%$ at the nominal $5\%$ threshold. In other words, the exact feasible score can be computed stably on genuine intraday data, but the SPY sample does not display persistent evidence for a supercritical regime.
Table 3: One-minute SPY boundary score: full-sample daily summary
Statistic Value
Trading days 2621
Mean $\widehat{T}_{n}^{\mathrm{m}}$ 0.005
Median $\widehat{T}_{n}^{\mathrm{m}}$ -0.185
Std. dev. $\widehat{T}_{n}^{\mathrm{m}}$ 0.916
Rejection share (10%) 10.4%
Rejection share (5%) 6.5%
Median $\widehat{\sigma}_{0}$ 0.0002
Table 4 breaks the sample into broad subperiods. The 2008–2009 block is somewhat more positive on average, while the 2020–2021 block is mildly negative, but none of the subperiod summaries suggests a sustained right-tail shift large enough to indicate repeated entry into $H>3/4$.
Table 4: One-minute SPY boundary score: subperiod summaries
Period Days Mean $T$ Median $T$ Reject 10% Reject 5% Median $\widehat{\sigma}_{0}$
2008-2009 385 0.118 -0.091 13.0% 7.3% 0.0002
2010-2019 1968 0.006 -0.188 10.3% 6.6% 0.0002
2020-2021 268 -0.158 -0.301 7.5% 4.9% 0.0002
Figure 3 plots the daily statistic together with its 60-day rolling mean. Crossings of the one sided critical lines occur, but they are not persistent and do not cumulate into a persistent upward regime. Figure 4 reports the 60-day rolling rejection share at the $5\%$ level. It fluctuates around the nominal benchmark and never points to a long interval of systematic supercritical rejection.
Figure 3: Daily exact feasible mfBm statistic on one-minute SPY returns, with 60-day rolling mean and one-sided critical lines.
Figure 4: 60-day rolling rejection share of the exact feasible mfBm statistic at the nominal $5\%$ level.
The empirical conclusion is therefore narrow but concrete. Once the section is built around the statistic developed earlier in the paper, the main message is not a recycled roughness estimate but a direct boundary statement: in this one-minute SPY sample, the exact feasible score does not deliver persistent evidence that the data live on the supercritical side of the semimartingale threshold.
This is the right interpretation of the empirical exercise. The object of interest in the present paper is not whether the data exhibit roughness in the now standard sense of very small Hurst exponents, but whether they cross the much higher boundary $H=3/4$ that separates the critical and supercritical regimes in mixed fractional models. In that sense, non-rejection in favor of $H>3/4$ should not be read as evidence against the broader rough-volatility literature. Rather, it says that the SPY sample stays on the critical or subcritical side of the critical boundary. Once that supercritical possibility is ruled out, the remaining question is how to estimate $H$ inside the subcritical region, and that is precisely the question treated by the roughness and fractional volatility literatures; see Gatheral et al. (2018), Bolko et al. (2023), Wang et al. (2023), and Shi et al. (2024).
9 Conclusion
This paper develops a boundary theory for high-frequency inference in mixed fractional models at the critical point $H=3/4$. Exactly there, the usual off-boundary normalizations fail when $H>3/4$, yet the transformed $(\sigma,H)$ block remains non-degenerate once the explicit linear component of the $H$-score has been removed. That single observation is enough to recover a workable local geometry and, from it, to derive exact critical score CLTs, explicit LAN expansions, and boundary-calibrated score tests for both the mfBm and mfOU experiments.
The broader message is that the critical boundary is statistically accessible. In a literature that has understandably concentrated on rough regimes with small Hurst exponents, our results address the complementary question of how to conduct disciplined inference when the object of interest lies near the semimartingale threshold itself. In that sense, the paper is meant to complement rather than compete with the rough-volatility literature represented by Bayer et al. (2016), Gatheral et al. (2018), Fukasawa (2021), and with the recent Journal of Econometrics contributions of Wang et al. (2023), Bolko et al. (2023), and Shi et al. (2024). Those papers show why fractional and rough dynamics matter empirically and how one estimates $H$ once the process is already understood to lie in a rough or otherwise subcritical region. The present one asks the prior question of whether a mixed fractional specification has crossed into the supercritical side of the boundary at all.
The score test analysis shows that the boundary experiment yields a feasible statistic with a clear nuisance orthogonal interpretation, and the Monte Carlo evidence shows why finite-sample calibration is delicate at the critical point. The intraday SPY illustration completes that picture on the data side. Once the exact feasible score is applied directly to regular-session one-minute returns, the empirical message is simple: the statistic is operational, but the sample does not exhibit persistent supercritical evidence. That is precisely the kind of statement the paper is designed to deliver. Therefore, the main remaining step is no longer about theory but about empirical scope: we need richer data sets and broader market panels to determine whether this largely subcritical pattern is unique to SPY or common across high-frequency equity benchmarks.
Appendix A Technical appendix: critical low frequency reductions and explicit constants
@@ A.4 Higher-order critical profiles and LAN remainders
$\left\lVert N_{\beta,n}^{\bullet}\right\rVert_{\mathrm{op}}=O\bigl((\log n)^{m(\beta)}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(N_{\beta,n}^{\bullet}\bigr)^{2}\Bigr)=O\bigl(n\Delta_{n}L_{n}^{2m(\beta)+1}\bigr),\qquad|\beta|=2,3.$ (56)
In particular, every additional $H$-derivative contributes at most one extra logarithmic factor at the operator level and two at the trace level.
@@
$\left\lVert\mathcal{R}_{2,n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}$ $=O\!\left(\frac{(\log n)^{2}}{n\Delta_{n}\,L_{n}^{3}}\right),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{2,n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)$ $=O\!\left((n\Delta_{n}L_{n})^{-1}\right),$ (58)
$\left\lVert\mathcal{R}_{2,n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}$ $=O\!\left(\frac{1+(\log n)^{2}/L_{n}^{3}}{n\Delta_{n}}\right),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{2,n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)$ $=O\!\left((n\Delta_{n})^{-1}\right),$ (59)
$\left\lVert\mathcal{R}_{3,n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}$ $=O\!\left(\frac{(\log n)^{3}}{(n\Delta_{n})^{3/2}L_{n}^{9/2}}\right),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{3,n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)$ $=O\!\left((n\Delta_{n})^{-2}L_{n}^{-2}\right),$ (60)
$\left\lVert\mathcal{R}_{3,n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}$ $=O\!\left(\frac{1+(\log n)^{3}/L_{n}^{9/2}}{(n\Delta_{n})^{3/2}}\right),$ $\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{3,n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)$ $=O\!\left((n\Delta_{n})^{-2}\right).$ (61)
@@
and proposition 20 shows that these three terms are respectively of operator orders
$O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),\qquad O\!\left(\frac{\log n}{n\Delta_{n}L_{n}^{2}}\right),\qquad O\!\left(\frac{(\log n)^{2}}{n\Delta_{n}L_{n}^{3}}\right),$
while all three have squared-trace order $O\bigl((n\Delta_{n}L_{n})^{-1}\bigr)$. This yields (58). For mfOU the additional $(\sigma,\alpha)$, $(H,\alpha)$ and $(\alpha,\alpha)$ terms have operator orders
$O\!\left(\frac{1}{n\Delta_{n}L_{n}^{1/2}}\right),\qquad O\!\left(\frac{\log n}{n\Delta_{n}L_{n}^{3/2}}\right),\qquad O\bigl((n\Delta_{n})^{-1}\bigr),$
and squared-trace orders
$O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),\qquad O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),\qquad O\bigl((n\Delta_{n})^{-1}\bigr),$
respectively; hence (59) follows.
For the cubic remainder, every monomial contains three local shifts and a third derivative. In mfBm the worst case is any monomial with total $m(\beta)=3$ after multiplying by $(\delta_{H,n})^{3}$, and all such contributions are bounded by
$(n\Delta_{n})^{-3/2}L_{n}^{-9/2}\cdot(\log n)^{3}$
@@ Appendix B Elementary logarithmic integrals
Data availability statement
The Monte Carlo evidence is generated from the designs described in Section 7. The empirical section in Section 8 uses the uploaded one-minute SPY data file 1_min_SPY_2008-2021.csv. The bundled files therefore contain the cleaned regular-session daily return panels, the exact feasible-score outputs, and the figures and tables reported in the empirical section. A fuller replication archive would add broader cross-asset intraday samples and parallel scripts for the mfOU side.